Hi...Can someone please help me with this problem:A 10-m ladder is leaning against a building. The bottom of the ladder is 5m from the building.How high is the top of the building?(To be rounded to the nearest tenths)..Thank you in advance!:)
Answer by mananth(16946)

The ladder , ground & the wall form a right triangle.
the ladder is the hypotenuse.
leg1^2+leg2^2=Hypotenuse^2
5^2+leg2^2=10^2
25+leg2^2=100
leg2^2=100-25
leg2^2=75
leg2 = sqrt75
leg 2= 8.7 feet the height of the wall